A block of mass m (initially at rest) is dropped from a height h landing onto a spring with spring/force constant k. Determine the maximum distance y that the spring will be compressed?


1
Expert's answer
2021-05-05T13:30:03-0400

"E_p=mgh,~E_s=\frac{ky^2}{2},"

"E_p=E_s,"

"y=\sqrt{\frac{2mgh}{k}}."